# Safe Text-to-Image Generation: Simply Sanitize the Prompt Embedding

### v1

@@ 7 Conclusion
In this paper, we investigate the risk of inappropriate generation in T2I models. To address this issue, we introduce Embedding Sanitizer (ES), a novel safe generation framework. ES successfully prevents potential inappropriate generation at the source by sanitizing the prompt embedding, thereby enhancing the safety and reliability of T2I model-generated content. Our framework not only effectively avoids generating inappropriate concepts but also provides improved interpretability and controllability. Moreover, ES’s modular design allows it to be compatible with existing safeguards, further enhancing the system’s flexibility. Through extensive comparisons across five datasets and with nine baseline methods, the results confirm that our approach achieves SOTA performance, effectively balancing robustness and fidelity.
@@ Appendix A Dataset Descriptions
  • Adversarial Prompt Datasets: Considering that certain safeguard baselines (e.g., SafeGen) are only applicable for erasing explicit content we focus on generating adversarial prompt datasets related to explicit content. (i) Ring-A-Bell (RAB). Based on the default settings from the Ring-A-Bell’s source code, we filter 931 prompts related to sexual content from the I2P dataset and apply the Ring-A-Bell attack to generate adversarial prompts, forming the Ring-A-Bell dataset. (ii) SneakyPrompt (SP). Similarly, we apply the SneakyPrompt attack on prompts in I2P related to sexual content, creating a SneakyPrompt dataset with 931 adversarial prompts. (iii) MMA-Diffusion (MMA). Additionally, we use a dataset of 1000 adversarial prompts published by MMA-Diffusion, focused on inducing T2I models to generate explicit content. These datasets allow us to evaluate ES ’s robustness under adversarial attacks.
  • COCO-2017: COCO-2017 is a widely used image-text dataset providing diverse scenes, object combinations, and real annotations, covering 80 object classes and background contexts from the real world. Following prior research, we use the validation set of COCO-2017 to evaluate ES ’s specificity and fidelity.
@@ Appendix B Baseline Descriptions
  • Concept Ablation (CA): We adopt Model-based CA to erase seven inappropriate concepts in I2P and fine-tuned the model’s cross-attention layer 1000 times at a learning rate of $2\times 10^{-7}$ according to the source code settings.

### v2

@@ 8 Conclusion
In this paper, we investigate the risks of unsafe generation in T2I models. To address this issue, we introduce a novel safe generation framework, Embedding Sanitizer (ES). ES enhances the safety and reliability of T2I outputs by sanitizing prompt embeddings, effectively preventing inappropriate generations at the source. Our framework not only mitigates the generation of harmful concepts but also improves interpretability and controllability. Moreover, the modular design of ES ensures compatibility with existing safety mechanisms, further enhancing system flexibility. Extensive comparisons across five datasets and eleven baseline methods demonstrate that our approach achieves SOTA performance, effectively balancing robustness and fidelity.
@@ Appendix A Dataset Descriptions
  • Adversarial Prompt Datasets: Considering that certain safeguard baselines (e.g., SafeGen and SC) are only applicable for erasing explicit content, we focus on generating adversarial prompt datasets related to explicit content. (i) Ring-A-Bell (RAB). Based on the default settings from the Ring-A-Bell’s source code, we filter 931 prompts related to sexual content from the I2P dataset and apply the Ring-A-Bell attack to generate adversarial prompts, forming the Ring-A-Bell dataset. (ii) SneakyPrompt (SP). Similarly, we apply the SneakyPrompt attack on prompts in I2P related to sexual content, creating a SneakyPrompt dataset with 931 adversarial prompts. (iii) MMA-Diffusion (MMA). Additionally, we use a dataset of 1000 adversarial prompts published by MMA-Diffusion, focused on inducing T2I models to generate explicit content. These datasets allow us to evaluate ES ’s robustness under adversarial attacks.
  • COCO-2017: COCO-2017 is a widely used benign image-text pair dataset that offers a diverse range of scenes. It contains images with various scenes and objects, along with their corresponding ground-truth text annotations, covering 80 object classes and real-world background contexts. Consistent with prior research [21, 8, 38], we use the validation set of COCO-2017 to evaluate ES ’s specificity and fidelity.
Figure 9: Comparison of the erasure rates for different exposed body parts between ES and safeguard baselines. SD-v1.4 generated a total of 3596 exposed body parts on the RAB dataset, while our method reduced this number to 31.
@@ Appendix B Baseline Descriptions
  • Prompt Optimizer For Safe T2I (POSI): POSI is a novel black-box approach for optimizing prompts in safe text-to-image generation. Following the default settings, we replicate the original method using LLaMA-7B with LoRA fine-tuning and GPT-3.5-generated toxic-clean prompt pairs. In the SFT phase, we set the learning rate to $5\times 10^{-5}$ and train for 3 epochs. During PPO, we use a learning rate of $1.9\times 10^{-5}$ and train for 1 epoch. The optimizer is trained to balance toxicity reduction with semantic preservation. We adopt the same prompt instruction as in the original work when using the model: ”You are an AI visual assistant, please help me to optimize the prompt used for text-to-image models to be non-toxic.”
@@
  • Rickrolling(RR): Rickrolling is a backdoor attack method targeting text-to-image synthesis models by subtly fine-tuning pre-trained text encoders. We adapt RR as a concept erasure technique by injecting backdoors into the CLIP text encoder. Twelve harmful concepts from the I2P dataset — including hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, and blood — are directly used as triggers and mapped to an empty target. The encoder is fine-tuned for 800 steps using SimilarityLoss, with a batch size of 128, loss weight of 0.1, and 8 data loader workers.
  • Moderator: is a policy-driven model editing framework for fine-grained content moderation in text-to-image models like Stable Diffusion. For each harmful concept in I2P, it defines a removal policy specifying the target object, editing method (removal), and related sub-concepts. To suppress a target concept, the model is fine-tuned for 600 steps with a batch size of 1 and a learning rate of 1e-5 to obtain a task vector capturing the concept direction. This task vector is then subtracted from the original model, effectively removing the undesired concept. Due to performance degradation when handling multiple concepts, Moderator is used exclusively for explicit content moderation.
Appendix C Evaluation of Erasure Effects on Six Exposed Body Parts
Figure 9 further demonstrates the defense effectiveness of the baseline methods and ES on the RAB adversarial prompt dataset. We use NudeNet[27] to detect exposed content in the generated images, which can identify various categories of exposure, such as female breasts (F_Breast) and male genitalia (M_Genitalia). It is noteworthy that the RAB attack generates more explicit content compared to I2P by adding extra pornographic semantics to the target prompt (the number of exposed content increases from 674 to 3596). The significant increase in explicit content enhances the output filter SC’s detection capability, resulting in SC achieving the best defense performance among the baselines under the RAB attack. In contrast, other baseline methods show weaker erasure effects in specific areas, such as Safe-CLIP’s limited effectiveness in erasing female breasts, and SafeGen’s poor erasure performance on the buttocks. However, ES effectively suppresses the generation of all exposed body parts, with an ER of 99.14% for total exposed areas, achieving SOTA performance.